# Solving Oscillator ODEs via Soft-constrained Physics-informed Neural Network with Small Data

### v1

@@ Abstract
This paper compared physics-informed neural network (PINN), conventional neural network (NN) and numerical discretization methods on solving differential equations through literature research and experimental validation. PINN inherits the strong expressive ability of NN and incorporates physical information and knowledge into neural network topology or learning architecture as well as computational process by means of hard-coding or soft constraints, which reduces data redundancy, improves generalization performance, and consequently increases computational efficiency, robustness, and interpretability. On the other hand, compared to numerical discretization methods (e.g., Runge-Kutta method, Finite Element Method), PINN has many advantages such as: 1) mesh-free, 2) seamlessly merging experimental data and tolerating noise, 3) effectively and efficiently evaluating new data samples, and 4) breaking the curse of dimensionality for high-dimensional, nonlinear, or inverse problems that are either not well solved or are expensive to solve by numerical discretization methods.
We formalized the mathematical framework and computational flow of the soft-constrained PINN method for solving differential equations (e.g., ODEs/PDEs). The working mechanism and its accuracy and efficiency were experimentally verified by solving typical linear and non-linear oscillator ODEs. We demonstrate that the implementation of the PINN method by DeepXDE is not only light code and efficient in training, but also flexible across CPU/GPU platforms. PINN greatly reduces the need for labeled data: when the nonlinearity of the ODE is weak, a very small amount of supervised training data plus a small amount of collocation points are sufficient to predict the solution; in the minimalist case, only one or two training points (including initial values) are needed for first- or second-order ODEs, respectively. Strongly nonlinear ODE also require only an appropriate increase in the number of training and collocation points, which still has significant advantages over conventional NN. We find that, with the aid of collocation points and the use of physical information, PINN has the ability to extrapolate data outside the time domain covered by the training set, and is robust to noisy data, thus with enhanced generalization capabilities. The training is accelerated when the gains obtained along with the reduction in the amount of data outweigh the delay caused by the increase in the loss function terms. The soft-constrained PINN method can easily impose a physical law (e.g., energy conservation) constraint by adding a regularization term to the total loss function, thus improving the solution performance of ODEs that obey this physical law. Furthermore, PINN can be used for stiff ODEs, PDEs, IDEs, PDEs, SDEs, and other types of differential equations, and is becoming a favorable catalyst for the era of Digital Twins.
Keywords: Physics-informed Neural Network(PINN); Oscillator ODEs; DeepXDE; Soft-constrained; Minimalistic Data; Noise; Energy Conservation Regularization; Non-linear
@@ 1.1 Why PINN: Where Neural Networks can be Improved
PINN incorporates physical information and knowledge into network topologies or computational processes as model priors, in a structured, modular neural network learning architecture. There are two specific implementations of PINN: the approaches of hard constraints and soft constraints. Hard constraints [LPY+21] generally ensure that physical knowledge or laws (e.g., differential equations, symmetries, conservation laws [MPS+19, WY23, YW24]), as well as boundary and initial conditions [LTSL20], are strictly adhered to by hard-coding or embedding them directly as part of the neural network architecture or computational process, which needs to be approximated by a specific network design or training strategy [WJX+22]. The soft constraints discussed in this paper, on the other hand, are indirectly achieved by adding data and governing equation residual or regularization terms to the loss function, which is easier to implement. This approach can be considered as a specific use case for multi-task learning. Soft constraints allow some flexibility or tolerance in satisfying physical laws or conditions, but require balancing predictive performance and physical accuracy by carefully adjusting the various weighting factors in the loss function.
Using these priors as residual or regularization terms allows the solution space of admissible solutions to the problem to be restricted to manageable limits. For example, in in-compressible fluid dynamics problems, any non-realistic flow solution that violates the law of mass conservation can be discarded [KKL+21]. In return, encoding this structured information into the learning algorithm amplifies the information of the data the algorithm sees, allowing it to quickly steer to the right solution [KKL+21]. The key property of PINN is that it generalizes well even with only a small amount of supervised or labeled training data. Furthermore, these prior knowledge or constraints can yield more interpretable learning methods. These methods remain robust to data imperfections (e.g., noise, missing or outliers) and provide accurate and physically consistent predictions that can even be used for extrapolation tasks.
@@ 1.2 PINN as a Complement or Alternative to Numerical Discretization
Differential equations such as Ordinary Differential Equations (ODEs) and Partial Differential Equations (PDEs) are important tools for expressing the laws of nature in science and engineering. Real-world physical systems can be modeled using these differential equations based on different domain-specific assumptions and simplifications. These models can be used to approximate the behavior of these systems [GKLS24]. The common representative numerical discretization methods for solving ODEs/PDEs are the Runge-Kutta method [HSH16] and the finite element method (FEM) [ALO+14], respectively. From a view of interpretability, these methods are well understood due to the simplicity of the underlying approximation - with readily available error estimates as well as convergence and stability guarantees.
@@
Great progress has been made in solving differential equations (ODE/PDE) by numerical discretization in order to simulate various types of field problems, and the theoretical foundations of this method are well established, with the advantages of high efficiency, high accuracy, and good stability. However, there are still some bottlenecks that severely limit its application [KKL+21]: (i) the curse of dimensionality problem, (ii) the mesh generation is still complicated, (iii) there are difficulties in merging experimental data, still unable to seamlessly incorporate noisy data into the existing algorithms, and (iv) it is not possible to solve the high-dimensional problems controlled by the parameterized PDEs, and so on. For example, an obvious disadvantage of FEM is that it relies on spatial discretization (spatial meshes plus large polynomial bases) and suffers from the curse of dimensionality: it is already difficult to use it in three dimensions, let alone for higher dimensional problems. Moreover, certain nonlinear and non-smooth partial differential equations are very difficult to discretize, e.g., due to general non-smooth behavior or singularities, which usually need to be solved on a very fine mesh. Since FEM is a mesh-based solver, solving over certain irregular domains requires customized methods that are challenging to design and solve [GKLS24]. In addition, solving inverse problems (e.g., for inferring material properties in functional materials or discovering missing physics in reaction transport) is often very expensive and requires complex formulas, new algorithms, and sophisticated computer codes. Moreover, solving realistic physical problems with missing, gapped, or noisy boundary conditions by conventional methods is currently impossible [KKL+21].
@@
Deep learning methods (e.g., PINN [RPK19], deep Ritz method [WY18], deep Galerkin method [SS18]) have achieved great success in solving high-dimensional problems and have recently become a common method for numerical solutions of various PDEs. They have the potential to overcome some of the challenges faced by the numerical discretization methods described above: (i) By utilizing automatic differentiation [BPRS17], the need for discretization is eliminated, giving the advantage of being mesh-free. (ii) Neural networks are able to represent more general functions than finite element bases, can break the curse of dimensionality [PMR+16], providing a promising direction for solving high-dimensional PDEs. (iii) Although training neural networks (the non-convex optimization problem) may become computationally intensive compared to numerical solvers such as FEM, it is very effective and efficient in evaluating new data samples (i.e., generalization ability). (iv) Its ability and flexibility to integrate data and knowledge provides a scalable framework for dealing with problems with imperfect knowledge or limited data. However, training deep neural networks requires a lot of data, which is not always available for related scientific problems, as mentioned earlier.
@@
Among them, PINN serves as a potentially disruptive technology that integrates physical information and knowledge into network topology or computational process. PINN can obey any symmetry, in-variance, or conservation principle derived, for example, from the laws of physics governing observational data, modeled by usual time-varying and nonlinear partial differential equations [RPK19]. This simple yet powerful construction allows us to address a wide range of problems in computer science, leading to the development of new data-efficient and physics-informed learning machines, new numerical solvers for ODEs/PDEs, and new data-driven methods for model inversion and system identification. It should be emphasized that PINN is not intended to replace FEM. PINN is particularly effective in solving, for example, the hypothetical and inverse problems described in this section; and for forward problems that do not require any data assimilation, existing grid-based numerical solvers currently outperform PINN [KKL+21, GKLS24]. Table 1 summarizes the advantages and disadvantages of the three types of methods (i.e., numerical discretization, conventional neural networks, and PINN) for solving differential equations.
@@ 1.3 Characteristics of the PINN Method
It is important to emphasize that, unlike any classical numerical method for solving differential equations (mesh-based or discrete step-size methods such as 4th order Runge-Kutta, FDM, FEM), the predictions of PINN are obtained without any discretization of the spatio-temporal domain and are able to break the curse of dimensionality [PMR+16, LMMK21].
PINN does not need to deal with prohibitively small step sizes, so it can easily handle irregular and moving domain problems [GKLS24] and scales well in higher dimensions. Automatic differentiation is a significant advantage of PINN over other classical numerical methods for solving DE, which utilizes the chain rule to back-propagate through the network and compute the derivatives, being able to accurately evaluate the differential operators at the configuration points with machine accuracy [BB23]. It does not depend on the discrete grid of the domain, but for which the sampling method becomes more important [WZT+23].
@@
Moreover, PINN [RPK19] can directly handle nonlinear problems (an inherent property of neural networks that are good at handling nonlinear problems) without any a priori assumptions, linearization, or local time stepping. PINN is effective and efficient in dealing with the ill-posed and inverse problems, e.g., forward and inverse problems with no initial or boundary conditions specified, or problems where some parameters in the PDEs are unknown [KKL+21], when classical numerical methods may fail.
@@
Method Numerical Discretization Neural Network PINN
@@
✓Advantages of conventional
Neural Networks +
@@
to overfitting,
Strong generalization
capability
✓Can handle imperfect data
@@ 1.4 Related Work & Our Contributions
Related Work on PINN Solving ODEs.  [BB23] carried out a systematic and comprehensive benchmarking of PINN solving typical linear/nonlinear oscillator ODEs based on PyTorch.  [Bat23b] addressed the problems of (i) limited known solution data, and (ii) integration intervals that, if too large, would make it difficult for the PINN to accurately predict the solution, when solving initial value problems (IVPs) for stiff ODEs. Improved strategies such as embedding more physical information, optimizing the training data loss to ensure that it considers all initial conditions more comprehensively, and incremental learning strategy, i.e., gradually increasing the integration intervals and optimizing the parameters of the PINN in each interval, and employing a shifting mesh to minimize the residuals of the differential equations, have resulted in the improved PINN having a higher accuracy and a more stable training in solving IVPs for stiff ODEs process. In addition, these improved strategies have been shown to be equally effective in solving boundary value problems (BVPs), such as the solution to the high Reynolds number steady-state convection-diffusion equation.  [Bat23a] further compared the performance of soft and hard constraints in solving higher-order Lane-Emden-Fowler type equations. The results show that soft constraints are more flexible and can handle many different physical scenarios and boundary conditions, but may require more training data and iterations to converge to a satisfactory solution. Hard constraints are able to satisfy the constraints of the differential equations more directly, and therefore may achieve higher accuracy in fewer iterations, but may require a more complex process of test solution selection and tuning.
@@
  • The advantages of PINN over conventional neural network, numerical discretization methods were surveyed in solving ODEs and also PDEs through relevant literature research (see Table 1).
@@
  • We formalized the mathematical framework and computational flow of the soft-constrained PINN method for solving differential equations (e.g., ODEs/PDEs). We then explored the working mechanism as well as the effectiveness and high efficiency of the PINN method experimentally, through solving typical oscillator (covering first and second order, linear and nonlinear) ODEs such as Primer, Van der Pol, and Duffing oscillators. Our validation shows that: 1) PINN embeds physical information and prior knowledge to mechanically reduce data redundancy and has the unique advantage of achieving strong generalization with small data. It greatly reduces the need for labeled data and can even predict solutions with minimalistic data (e.g., linear first- and second-order differential equations require only one and two training data containing initial value, plus a few collocation points, respectively), and also improves computational efficiency. 2) PINN is robust to noisy data and provides accurate and physically consistent predictions and even extrapolates outside the time domain of the training set, the generalization ability is enhanced. And with 3) better intepretability.
@@
  • Taking the convergence of a second-order nonlinear Duffing oscillator with minimalistic data (2 training points containing initial value plus a small number of collocation points) as a validation example, it is shown that the soft-constrained PINN method can easily impose a physical law (e.g., energy conservation) constraint by adding a regularization term to the total loss function, thus improving the solution performance of ODEs that obey this physical law.
@@
  • The advantages of the DeepXDE-based implementation of the PINN method, such as light code and fast training speed, are demonstrated.

### v2

# Solving Oscillator Ordinary Differential Equations via Soft-constrained Physics-informed Neural Network with Small Data

@@ Abstract
This paper compared physics-informed neural network (PINN), conventional neural network (NN) and traditional numerical discretization methods on solving differential equations (DEs) through literature investigation and experimental validation. We focused on the soft-constrained PINN approach and formalized its mathematical framework and computational flow for solving Ordinary DEs and Partial DEs (ODEs/PDEs). The working mechanism and its accuracy and efficiency were experimentally verified by solving typical linear and non-linear (e.g., Primer, Van der Pol, Duffing) oscillator ODEs. We demonstrate that the DeepXDE-based implementation of PINN is not only light code and efficient in training, but also flexible across CPU/GPU platforms. PINN greatly reduces the need for labeled data: when the nonlinearity of the ODE is weak, a very small amount of supervised training data plus a few unsupervised collocation points are sufficient to predict the solution; in the minimalist case, only one or two training points (with initial values) are needed for first- or second-order ODEs, respectively. Strongly nonlinear ODE also require only an appropriate increase in the number of training points, which still has significant advantages over conventional NN. We also find that, with the aid of collocation points and the use of physical information, PINN has the ability to extrapolate data outside the time domain of the training set, and especially is robust to noisy data, thus with enhanced generalization capabilities. Training is accelerated when the gains obtained along with the reduction in the amount of data outweigh the delay caused by the increase in the loss function terms. The soft-constrained PINN can easily impose a physical law (e.g., conservation of energy) constraint by adding a regularization term to the total loss function, thus improving the solution performance to ODEs that obey this physical law. Furthermore, PINN can also be used for stiff ODEs, PDEs, and other types of DEs, and is becoming a favorable catalyst for the era of Digital Twins.
Keywords: PINN; Oscillator ODEs; DeepXDE; Soft-constrained; Minimalistic Data; Noise; Regularization of Conservation of Energy; Non-linear
@@ 1.1 Why PINN: Where Neural Networks can be Improved
PINN incorporates physical information and knowledge into network topology or computational processes as model priors, in a structured, modular neural network learning architecture. There are two specific implementations of PINN: the approaches of hard constraints and soft constraints. Hard constraints [LPY+21] generally ensure that physical knowledge or laws (e.g., differential equations, symmetries, conservation laws [MPS+19, WY23, YW24]), as well as boundary and initial conditions [LTSL20], are strictly adhered to by hard-coding or embedding them directly as part of the neural network architecture or computational process, which needs to be approximated by a specific network design or training strategy [WJX+22]. The soft constraints discussed in this paper, on the other hand, are indirectly achieved by adding data and governing equation residual or regularization terms to the loss function, which is easier to implement. This approach can be considered as a specific use case for multi-task learning. Soft constraints allow some flexibility or tolerance in satisfying physical laws or conditions, but require balancing predictive performance and physical accuracy by carefully adjusting the various weighting factors in the loss function.
Using these priors as residual or regularization terms allows the solution space of admissible solutions to the problem to be restricted to manageable limits. For example, in in-compressible fluid dynamics problems, any non-realistic flow solution that violates the law of mass conservation can be discarded. In return, encoding this structured information into the learning algorithm amplifies the information of the data the algorithm sees, allowing it to quickly steer to the right solution [KKL+21]. The key property of PINN is that it generalizes well even with only a small amount of supervised or labeled training data. Furthermore, these prior knowledge or constraints can yield more interpretable learning methods. These methods remain robust to data imperfections (e.g., noise, missing or outliers) and provide accurate and physically consistent predictions that can even be used for extrapolation tasks [KKL+21].
@@ 1.2 PINN as a Complement or Alternative to Numerical Discretization
Differential equations such as ordinary differential equations (ODEs) and partial differential equations (PDEs) are important tools for expressing the laws of nature in science and engineering. Real-world physical systems can be modeled using these differential equations based on different domain-specific assumptions and simplifications. These models can be used to approximate the behavior of these systems [GKLS24]. The common representative numerical discretization methods for solving ODEs/PDEs are the Runge-Kutta method [HSH16] and the finite element method (FEM) [ALO+14], respectively. From a view of interpretability, these methods are well understood due to the simplicity of the underlying approximation, with readily available error estimates as well as convergence and stability guarantees.
@@
Great progress has been made in solving differential equations (ODE/PDE) by numerical discretization in order to simulate various types of field problems, and its theoretical foundations are well established, with the advantages of high efficiency, high accuracy, and good stability. However, there are still some bottlenecks that severely limit its application [KKL+21]: (i) the curse of dimensionality problem, (ii) the mesh generation is still complicated, (iii) there are difficulties in merging experimental data, still unable to seamlessly incorporate noisy data into the existing algorithms, and (iv) it is not possible to solve the high-dimensional problems controlled by the parameterized PDEs, and so on. For example, an obvious disadvantage of FEM is that it relies on spatial discretization (spatial meshes plus large polynomial bases) and suffers from the curse of dimensionality: it is already difficult to use it in three dimensions, let alone for higher dimensional problems. Moreover, certain nonlinear and non-smooth PDEs are very difficult to discretize, e.g., due to general non-smooth behavior or singularities, which usually need to be solved on a very fine mesh. Since FEM is a mesh-based solver, solving over certain irregular domains requires customized methods that are challenging to design and solve [GKLS24]. In addition, solving inverse problems (e.g., for inferring material properties in functional materials or discovering missing physics in reaction transport) is often very expensive and requires complex formulas, new algorithms, and sophisticated computer codes. Moreover, solving realistic physical problems with missing, gapped, or noisy boundary conditions by conventional methods is currently impossible [KKL+21].
@@
Deep learning methods (e.g., PINN [RPK19], deep Ritz method [WY18], deep Galerkin method [SS18]) have achieved great success in solving high-dimensional problems and have recently become a common method for numerical solutions of various PDEs. They have the potential to overcome some of the challenges faced by the numerical discretization methods described above: (i) By utilizing automatic differentiation [BPRS17], the need for discretization is eliminated, giving the advantage of being mesh-free. (ii) Neural networks are able to represent more general functions than finite element bases, can break the curse of dimensionality [PMR+16], providing a promising direction for solving high-dimensional PDEs. (iii) Although training neural networks (the non-convex optimization problem) may become computationally intensive compared to numerical solvers such as FEM, it is very effective and efficient in evaluating new data samples. (iv) Its ability and flexibility to integrate data and knowledge provides a scale-able framework for dealing with problems with imperfect knowledge or limited data. However, training deep neural networks requires a lot of data, which is not always available for related scientific problems, as mentioned earlier.
@@
Among them, PINN serves as a potentially disruptive technology that integrates physical information and knowledge into network topology or computational process. PINN can obey any symmetry, in-variance, or conservation principle derived, for example, from the laws of physics governing observational data, modeled by usual time-varying and nonlinear PDEs [RPK19]. This simple yet powerful construction allows us to address a wide range of problems in computer science, leading to the development of new data-efficient and physics-informed learning machines, new numerical solvers for ODEs/PDEs, and new data-driven methods for model inversion and system identification. It should be emphasized that PINN is not intended to replace FEM. PINN is particularly effective in solving, for example, the hypothetical and inverse problems described in this section; and for forward problems that do not require any data assimilation, existing grid-based numerical solvers currently outperform PINN [KKL+21, GKLS24]. Table 1 summarizes the advantages and disadvantages of the three types of methods, i.e., numerical discretization, conventional neural networks (NN), and PINN for solving differential equations.
@@ 1.3 Characteristics of the PINN Method
It is important to emphasize that, unlike any classical numerical method for solving differential equations (mesh-based or discrete step-size methods such as 4th-order Runge-Kutta, FDM, FEM), the predictions of PINN are obtained without any discretization of the spatio-temporal domain and it is able to break the curse of dimensionality [PMR+16, LMMK21]. PINN does not need to deal with prohibitively small step sizes, so it can easily handle irregular and moving domain problems [GKLS24] and scales well in higher dimensions. Automatic differentiation is a significant advantage of PINN over other classical numerical methods for solving DEs, which utilizes the chain rule to back-propagate through the network and compute the derivatives, being able to accurately evaluate the differential operators at the configuration points with machine accuracy [BB23]. It does not depend on the discrete grid of the domain, but for which the sampling method becomes more important [WZT+23, YKH+24].
@@
Moreover, PINN can directly handle nonlinear problems (an inherent property of neural networks that are good at handling nonlinear problems) without a priori assumptions, linearization, or local time stepping [RPK19]. PINN is effective and efficient in dealing with the ill-posed and inverse problems, e.g., forward and inverse problems with no initial or boundary conditions specified, or problems where some parameters in the PDEs are unknown [KKL+21], when classical numerical methods may fail.
@@
Method Numerical Discretization Conventional NN PINN
@@
✓Advantages of
conventional NN +
@@
to overfitting, strong
generalization capability
✓Robust to imperfect data
@@ 1.4 Related Work & Our Contributions
Related Work on PINN Solving Oscillator ODEs.  [BB23] carried out a systematic and comprehensive benchmarking of PINN solving typical linear/nonlinear oscillator ODEs based on PyTorch.  [Bat23b] addressed the problems of (i) limited known solution data, and (ii) integration intervals that, if too large, would make it difficult for the PINN to accurately predict the solution, when solving initial value problems (IVPs) for stiff ODEs. Improved strategies such as embedding more physical information, optimizing the training data loss to ensure that it considers all initial conditions more comprehensively, and incremental learning strategy, i.e., gradually increasing the integration intervals and optimizing the parameters of the PINN in each interval, and employing a shifting mesh to minimize the residuals of the DEs, have resulted in the improved PINN having a higher accuracy and a more stable training in solving IVPs for stiff ODEs process. In addition, these improved strategies have been shown to be equally effective in solving boundary value problems (BVPs), such as the solution to the high Reynolds number steady-state convection-diffusion equation.  [Bat23a] further compared the performance of soft and hard constraints in solving higher-order Lane-Emden-Fowler type equations. The results show that soft constraints are more flexible and can handle many different physical scenarios and boundary conditions, but may require more training data and iterations to converge to a satisfactory solution. Hard constraints are able to satisfy the constraints of the differential equations more directly, and therefore may achieve higher accuracy in fewer iterations, but may require a more complex process of test solution selection and tuning.
@@
  • The advantages of PINN over conventional neural networks, traditional numerical discretization methods were surveyed and summarized in solving ODEs and also PDEs through relevant literature research (see Table 1).
@@
  • We formalized the mathematical framework and computational flow of the soft-constrained PINN approach to solving ODEs/PDEs. We then explored the working mechanism as well as the effectiveness and high efficiency of the PINN method experimentally, through solving typical oscillator (covering first- and second-order, linear and nonlinear) ODEs such as Primer, Van der Pol, and Duffing oscillators. Our verification shows that: 1) PINN embeds physical information and prior knowledge to mechanically reduce data redundancy and has the unique advantage of achieving strong generalization with small data. It greatly reduces the need for labeled data and can even predict solutions with minimalistic data (e.g., linear first- and second-order differential equations require only one and two training data containing initial values, plus a few collocation points, respectively), and also improves computational efficiency. 2) PINN is robust to noisy data and provides accurate and physically consistent predictions and even extrapolates outside the time domain of the training set, therefore the generalization ability is enhanced. And with 3) better intepretability.
@@
  • Taking the convergence of a second-order nonlinear Duffing oscillator with minimalistic data (2 training points containing initial values plus some collocation points) as a validation example, it is shown that the soft-constrained PINN can easily impose a physical law (e.g., conservation of energy) constraint by adding a regularization term to the total loss function, thus improving the solution performance to ODEs that obey this physical law.
@@
  • The advantages of the DeepXDE-based implementation of the PINN method, such as light code, fast training speed, and flexible across CPU/GPU platforms were demonstrated.
This paper is organized as follows. In 1 Introduction, we point out that the success of deep neural networks relies on big data and strong computational power, and that there are improvements to be made especially in small data scenarios, leading to the PINN method that can leverage prior information and knowledge (1.1). In terms of solving ODEs/PDEs, it is then adequately compared with conventional neural networks and traditional numerical discretisation methods (1.2), thereby distilling the features and advantages of the PINN method (1.3). Then, using solving Oscillator ODEs as an analytical and experimental case study, we briefly describe the related work and summarise the contributions of ours (1.4). In 2 Method & Result, 2.1 establishes the mathematical framework and computational flow of the soft-constraint PINN approach focused in this paper and the DeepXDE implementation. 2.2 demonstrates the working mechanism of the soft-constraint PINN approach and its accuracy and efficiency in details with examples of typical oscillator ODEs (covering first-/second- order, linear/nonlinear): 1) PINN greatly reduces the need for labeled data, works even with minimalistic data, speeds up training, and can also predict data outside the time domain of the training set (2.2.1). 2) The nonlinear case is consistent with the linear case (just need to add extra training points), and PINN has good robustness to noisy data, reflecting its strong generalization performance. DeepXDE-based implementation of PINN is light code with high training efficiency, and flexible across GPU/CPU platforms; the acceleration effect of GPU is obvious, however, the small data characteristics of PINN make it also very efficient to train on CPU (2.2.2). In particular, the comparison of convergence results of second-order nonlinear Duffing oscillator with or without the regularization of conservation of energy under minimalistic data demonstrates that soft-constrained PINN can easily integrate the physical laws to improve the solution results (2.2.3). Finally, in 3 Concluding Remarks, this work is summarized in 3.1 and 3.2 outlooks some future steps.